Prove that (1-w+w2)(1+w-w2)=4 if 1,w,w2 is the cube unit

Grade:11

1 Answers

Arun
6 years ago
we know 1+w+w2 = 0
hence we can write 1 + w = –w2 & 1 + w2 = –w
now ( –w –w)( –w2 –w2) = (–2w)* (–2w2)
4 w3 = 4  hence proved(we know that w3 = 1)